# Learning an Adaptive and View-Invariant Vision Transformer for Real-Time UAV Tracking

### v2

@@ II-C Multi-Teacher Knowledge Distillation
Multi-teacher knowledge distillation (MD) leverages knowledge from multiple off-the-shelf pre-trained teacher models to guide the student, enhancing its generalization by inheriting diverse knowledge from various teachers [24, 25]. It has primarily been studied in the contexts of image classification [68, 69, 70, 27], action recognition [71] and visual retrieval [28]. For instance, [27] proposed the multi-teacher distillation method for class incremental learning, using weight permutation, feature perturbation, and diversity regularization to ensure diverse mechanisms in teachers. To better leverage the complementary knowledge of each modality, [70] propose a multi-teacher multi-modal knowledge distillation framework to guide the training of the multi-modal fusion network and further improve the multi-modal feature fusion process. Whiten-MTD [28] aligns the outputs of teacher models by whitening their similarity distributions and identifies the most effective fusion strategy for their multi-teacher distillation framework through empirical analysis. To the best of our knowledge, existing knowledge distillation methods for visual tracking rely on a single-teacher framework [72, 73], limiting their ability to fully utilize the diverse knowledge from multiple teachers, potentially leading to suboptimal tracking performance. In this work, we propose a simple yet effective MI maximization-based multi-teacher knowledge distillation framework, integrated into AVTrack, to develop a more efficient UAV tracker. For the teacher models, we utilize all three off-the-shelf trackers from AVTrack (i.e., AVTrack-DeiT, AVTrack-ViT, and AVTrack-EVA), offering diverse and high-quality teachers without requiring additional operations, such as those in [27], to implement varied mechanisms in the teachers. To ensure the student model captures the most relevant information from the teacher models’ representations, we propose maximizing the MI between the averaged softened feature representation of the multi-teacher models and the student model’s softened feature representation. Averaging the predictions of all teacher models is a common and effective method, as it reduces biases and provides a more objective output than any single teacher’s prediction [74, 75, 76, 77].
@@ III-A Overview
The proposed AVTrack introduces a novel single-stream tracking framework, featuring an adaptive ViT-based backbone and a prediction head. To improve the efficiency of the ViT, each transformer block in the ViT-based backbone, except for the first $n_{f}$ blocks, incorporates an activation module (AM). This module is trained to adaptively determine whether to activate the associated transformer block. The framework takes a pair of images as input, comprising a template denoted as $Z\in\mathbb{R}^{3\times H_{z}\times W_{z}}$ and a search image denoted as $X\in\mathbb{R}^{3\times H_{x}\times W_{x}}$. These images are split into patches of size $P\times P$, and the number of patches for $Z$ and $X$ are $P_{z}=H_{z}\times W_{z}/P^{2}$ and $P_{x}=H_{x}\times W_{x}/P^{2}$, respectively. The features extracted from the backbone are fed into the prediction head to generate tracking results. To obtain view-invariant representation with ViTs, we maximize the mutual information (MI) between the feature representations of two different views of the target, i.e., the template image and the target patch in the search image. During the training phase, since the ground truth localization of the target in the search image is known, we can obtain the feature representation of the subsequent view from the representation of the search image using interpolation techniques.
Building on the proposed AVTrack, we introduce a simple yet effective MI maximization-based multi-teacher knowledge distillation (MD) framework to develop a more efficient UAV tracker, called AVTrack-MD. During the training of the AVTrack-MD model (i.e., the student models), the AVTrack’s (i.e., the teacher models) weights remain fixed while both the teacher and student models receive inputs $Z$ and $X$. Let $\mathfrak{B}_{T}$ and $\mathfrak{B}_{S}$ represent the backbones of the teacher and student, respectively. In our implementation, $\mathfrak{B}_{S}$ shares the same structure as the $\mathfrak{B}_{T}$ but uses a smaller ViT layer. Feature-based knowledge distillation transfers knowledge from the teacher models’ backbone features to the student model by maximizing the MI between the averaged softened features of the multi-teacher models and the student model’s softened features.
The details of the these design principles will be elaborated in the subsequent subsections.
@@ III-B Activation Module (AM)
The Activation Module selectively activates transformer blocks based on input, enabling adaptive adjustments to the ViT’s architecture and ensuring that only the essential blocks are activated for effective tracking. To enhance efficiency, the AM processes a subset of tokens representing both the target template and the search image. Its output determines whether the current Transformer block is activated; if not, the tokens from the preceding block are bypassed. Specifically, let’s consider the $i$-th layer ($i>n_{f}$). We denote the total number of tokens by $\mathcal{K}$, the embedding dimension of each token by $d$, and all the tokens output by the $(i-1)$-th layer by $\mathbf{t}_{1:\mathcal{K}}^{i-1}(Z,X)\in\mathbb{R}^{\mathcal{K}\times d}$. The slice of all tokens generated by the $(i-1)$-th Transformer block is expressed as $\mathbf{e}_{1}^{\textup{T}}\mathbf{t}_{1:\mathcal{K}}^{i-1}(Z,X):=\textbf{r}^{i-1}\in\mathbb{R}^{\mathcal{K}}$, where $\mathbf{e}_{1}^{\textup{T}}=[1,0,...,0]\in\mathbb{R}^{\mathcal{K}}$ is a standard unit vector in $\mathbb{R}^{\mathcal{K}}$, the linear layer is denoted by $\mathfrak{L}^{i}$. Formally, the Activation Module (AM) at layer $i$ is expressed as:
@@
where $p^{i}\in[0,1]$ represents the activation probability of the $i$-th Transformer block, $\sigma(x)=1/(1+e^{-x})$ indicates the sigmoid function. The sigmoid function is versatile and finds applications in various machine learning tasks where non-linear transformations or probabilistic outputs are required. In our work, the activation mechanism in each Transformer block is determined by the sigmoid function, which helps introduce non-linearity into the activation module’s decision-making process. Specifically, the sigmoid function maps the output value of the linear layer to a range between 0 and 1, which represents the activation probability of the associate Transformer block. If $p^{i}>\beta$, where $\beta\in(0.5,1)$ is the activation probability threshold, the transformer block at layer $i$ will be activated; otherwise, it is deactivated and the output tokens from the $(i-1)$-th layer will be fed into the $(i+1)$-th block directly. Let $\mathcal{N}$ denote the total number of transformer blocks in the given ViT. Theoretically, deactivating all $\mathcal{N}$ blocks simultaneously would result in no correlation being computed between the template and search image. To avoid such unfavorable conditions, the first $n_{f}$ layers are mandated to remain activated. This strategy helps alleviate computational burdens associated with AM, as these initial layers are typically essential, providing foundational information on which high-level and more abstract features and representations can be built. Another extreme case occurs when all Transformer blocks are activated for any input, allowing the model to more efficiently minimize classification and regression losses, as larger models have greater fitting capacity. To address this, we introduce a block sparsity loss, $\mathcal{L}_{spar}$, which penalizes a higher mean probability across all adaptive layers, encouraging the deactivation of many blocks on average to improve efficiency. The block sparsity loss is defined as follows:
@@ III-C View-Invariant Representations (VIR) via Mutual Information Maximization
Figure 3: An overview of the proposed MI maximization-based multi-teacher knowledge distillation. Note that $\mathcal{F}_{\tau}’$ ’ denotes the randomly shuffled of $\mathcal{F}_{\tau}$.
@@ III-D MI Maximization-based Multi-Teacher Knowledge Distillation (MD)
In distillation training, the student model is trained using a weighted sum of $\mathcal{L}_{MD}$ and the total loss function used for training the teacher model. $\tau$
@@ III-E Prediction Head and Training Objective
For the tracking task, we employ the weighted focal loss [83] for classification and a combination of $L_{1}$ loss and GIoU loss [84] for bounding box regression. The overall loss function is formulated as:
$\mathcal{L}_{overall}=\mathcal{L}_{cls}+\lambda_{iou}\mathcal{L}_{iou}+\lambda_{L_{1}}\mathcal{L}_{L1}+\gamma\mathcal{L}_{spar}+\kappa\mathcal{L}_{vir}$ (9)
where the constants $\lambda_{iou}=2$ and $\lambda_{L_{1}}=5$ are the same as in [12], $\gamma$ and $\kappa$ are set to $50$ and $0.0001$, respectively. Our framework is trained end-to-end with the overall loss $\mathcal{L}_{overall}$ after the pretrained weights of the ViT for image classification is loaded. After this training, we apply the proposed MI maximization-based multi-teacher knowledge distillation framework to obtain a student model that better balances performance and efficiency. Specifically, during the distillation phase, the distillation loss $\mathcal{L}_{MD}$ is added to the overall loss from the previous training stage, yielding the total loss $\mathcal{L}_{overall}^{*}$ for end-to-end distillation training as follows:
$\mathcal{L}_{overall}^{*}=\mathcal{L}_{overall}+\eta\mathcal{L}_{MD}$ (10)
where the weight coefficients in $\mathcal{L}_{overall}$ remain the same as those used during teacher model training, and the weight $\eta$ of $\mathcal{L}_{MD}$ is set to $0.1\times 10^{-3}$.
@@ IV Experiments
In this section, we evaluate our tracker on six publicly widely-used UAV tracking benchmarks, including DTB70 [85], UAVDT [86], VisDrone2018 [87], UAV123 [88], UAV123@10fps [88], and WebUAV-3M [105]. Specifically, we conduct a thorough comparison with over 38 existing state-of-the-art (SOTA) trackers, using their results obtained from running the official codes with the corresponding hyperparameters. Please note that our evaluation is performed on a PC that was equipped with an i9-10850K processor (3.6GHz), 16GB of RAM, and an NVIDIA TitanX GPU. To provide a clearer evaluation comparison with different types of trackers, we categorize them into two groups: lightweight trackers and deep trackers.
@@ IV-A Implementation Details
Model. The AVTrack utilizes three different lightweight ViTs as backbones to build three trackers for evaluation: AVTrack-ViT, AVTrack-DeiT, and AVTrack-EVA. Building on it, AVTrack-MD introduces the novel MI maximization-based multi-teacher knowledge distillation (MD) framework to develop three corresponding improved trackers: AVTrack-MD-ViT, AVTrack-MD-DeiT, and AVTrack-MD-EVA. The head and input sizes for AVTrack and AVTrack-MD are configured identically. The head consists of a stack of four Conv-BN-ReLU layers, with input sizes set to $256^{2}$ for the search image and $128^{2}$ for the template image.
@@ IV-B Comparison with Lightweight Trackers
1) Overall Performance: We compare our methods with 22 SOTA lightweight trackers, inclunding KCF [89], fDSST [90], ECO_HC [91], MCCT_H [92], STRCF [93], ARCF [33], AutoTrack [11], RACF [94], HiFT [34], SiamAPN [95], P-SiamFC++ [36], TCTrack [35], TCTrack++ [96], SGDViT [97], ABDNet [98], DRCI [99], PRL-Track [100], Aba-ViTrack [7], LiteTrack [101], SMAT [102], LightFC [103], and SuperSBT [104], across five renowned UAV tracking datasets: DTB70 [85], UAVDT [86], VisDrone2018 [87], UAV123 [88], and UAV123@10fps [88]. The evaluation results are presented in Table I. From table, our AVTrack-DeiT outperforms all SOTA trackers except Aba-ViTrack [7] across these five benchmarks in terms of average (Avg.) precision (Prec.) and success rate (Succ.). Specifically, RACF [94] exhibits the best performance among DCF-based trackers, with Avg. Prec. and Succ. of 74.6% and 51.2%, respectively. Among CNN-based trackers, DRCI [99] stands out with the highest Avg. Prec. and Succ. of 79.8% and 59.1%. However, even the best methods among DCF- and CNN-based trackers fall short of the worst method among ViT-based trackers. All ViT-based trackers achieve Avg. Prec. and Succ. exceeding 80.0% and 62.0%, respectively, clearly surpassing the CNN-based methods and significantly outperforming the DCF-based ones. Regarding GPU speed, the top three trackers are our proposed methods: AVTrack-MD-EVA, AVTrack-MD-DeiT, and AVTrack-MD-ViT, achieving tracking speeds of 334.4 FPS, 310.6 FPS, and 303.1 FPS, respectively. In terms of CPU speed, all our methods deliver real-time performance on a single CPU [^1] It is important to note that the real-time performance discussed in this work is applicable only to platforms that are similar to or more advanced than ours.. Obviously, DCF-based trackers are the most efficient UAV trackers, as all methods achieving speeds above 80 FPS belong to this category. However, these fastest DCF-based trackers are unable to achieve performance comparable to our methods. For example, the fastest KCF [89] tracker attains only 31.6% in average success rate, which is about half that of our methods. Additionally, DCF-based methods often incur substantial costs to enhance tracking precision. Specifically, RACF exhibits the best performance among DCF-based trackers; however, it runs at only 35.6 FPS, which is significantly lower than the slowest CPU speed of our trackers. Although CNN-based trackers can achieve tracking speeds comparable to our ViT-based trackers, the latter significantly outperform the former in overall performance. Although Aba-ViTrack achieves the highest average precision of 85.3% and the highest average success rate of 64.7%., our AVTrack-DeiT secures second place with only slight gaps of 1.2% and 0.3%, respectively. Furthermore, our AVTrack-DeiT demonstrates higher speeds than Aba-ViTrack. Notably, AVTrack-DeiT runs up to 41%/18% faster than Aba-ViTrack in terms of GPU/CPU speed. When the proposed MI-based multi-teacher knowledge distillation framework (MD) is applied to AVTrack-*, the resulting student model, AVTrack-MD-*, demonstrates a smaller performance drop, with some even achieving performance improvements, all while delivering a significant speed boost. Specifically, AVTrack-MD-* models show varying trade-offs: AVTrack-MD-ViT achieves 0.5% and 0.9% gains in Avg. Prec. and Succ., respectively, with 21.1% GPU and 8.7% CPU speed improvements; AVTrack-MD-EVA gains 0.4% in precision and a minor 0.1% drop in success rate, while also improving GPU and CPU speeds by 17.9% and 6.8%; AVTrack-MD-DeiT has a slight performance drop of 0.5% in precision and 0.3% in success rate but still benefits from significant 20.9% GPU and 8.9% CPU speed gains. These results highlight the advantages of our method and justify its SOTA performance for UAV tracking.
@@
PRL-Track 62.3 47.1 12.1 7.4 33.8
LiteTrack 69.1 54.1 28.3 7.3 32.1
SMAT 68.9 53.9 8.6 3.2 32.7
Aba-ViTrack 70.4 55.3 8.0 2.4 37.3
ABDNet 63.9 48.7 12.3 8.3 33.2
SGDViT 61.3 45.7 23.3 11.3 31.7
TCTrack++ 63.9 48.3 17.6 8.8 32.5
TCTrack 61.9 45.7 8.5 6.9 34.4
SiamAPN 62.5 45.1 14.5 7.9 35.2
HiFT 60.9 45.5 9.9 7.2 35.6
@@
2) Performance on WebUAV-3M [105]: As illustrated in Table II, we have conducted a comparison of our AVTrack-DeiT and AVTrack-MD-DeiT with ten lightweight SOTA trackers on WebUAV-3M. Our AVTrack-DeiT surpasses all other lightweight trackers in success rate, achieving a 1.1% improvement over the second-place Aba-ViTrack while maintaining a comparable precision with only a 0.4% difference. Notably, our AVTrack-MD-DeiT experiences less than a 1.0% drop in both precision and success rate, while improving AGX speed by 9.2%. These results further underscore the effectiveness of our approach.
@@
3) Efficiency Comparison: To further demonstrate that our tracker achieves a better trade-off between accuracy and efficiency, we compare our methods against ten SOTA lightweight trackers, in terms of floating point operations (FLOPs), the number of parameters (Params.), and the speed (AGX.FPS) on the Nvidia Jetson AGX Xavier edge device. The results are shown in Table II. Notably, since our AVTrack-DeiT tracker features adaptive architectures, the FLOPs and parameters range from minimum to maximum values. As observed, both the minimum FLOPS and Params of our AVTrack-DeiT are notably lower than those of all the SOTA trackers, and even its maximum values are lower than those of most SOTA trackers. Furthermore, our AVTrack-MD-DeiT achieves the lowest FLOPs (1.5G) and parameters (5.3M), except for AVTrack-DeiT, while also delivering the fastest speed at 46.1 AGX.FPS. This comparison in terms of computational complexity also underscores the efficiency of our methods.

### v3

@@ II-C Multi-Teacher Knowledge Distillation
Multi-teacher knowledge distillation (MD) leverages knowledge from multiple off-the-shelf pre-trained teacher models to guide the student, enhancing its generalization by inheriting diverse knowledge from various teachers [36, 37]. It has primarily been studied in the contexts of image classification [69, 70, 39], action recognition [71] and visual retrieval [40]. For instance, MTD [39] explored a multi-teacher distillation approach for class-incremental learning, leveraging weight permutation, feature perturbation, and diversity regularization to ensure diverse mechanisms in teachers. To better leverage the complementary knowledge of each modality, M2KD [70] propose a multi-teacher multi-modal knowledge distillation framework to guide the training of the multi-modal fusion network and further improve the multi-modal feature fusion process. Whiten-MTD [40] aligns the outputs of teacher models by whitening their similarity distributions and identifies the most effective fusion strategy for their multi-teacher distillation framework through empirical analysis. To the best of our knowledge, existing knowledge distillation methods for visual tracking rely on a single-teacher framework [72, 73], limiting their ability to fully utilize the diverse knowledge from multiple teachers, potentially leading to suboptimal tracking performance. In this work, we propose a simple yet effective MI maximization-based multi-teacher knowledge distillation framework, integrated into AVTrack, to develop a more efficient UAV tracker. We adopt all three off-the-shelf trackers from AVTrack (i.e.,AVTrack-DeiT, AVTrack-ViT, and AVTrack-EVA) as teacher models, which offer a diverse and high-quality selection of teachers, eliminating the need for additional operations, like those in [39], to implement varied mechanisms in the teachers. To ensure the student model captures the most relevant information from the teacher models’ representations, we propose maximizing the MI between the averaged softened feature representation of the multi-teacher models and the student model’s softened feature representation. Averaging the predictions of all teacher models is a common and effective method, as it reduces biases and provides a more objective output than any single teacher’s prediction [74, 75, 76, 77].
@@ III Method
Figure 3: An overview of the AM’s internal working mechanism. Note that $p$ and $\beta$ denote the activation probability and its threshold, respectively.
@@ III-A Overview
The proposed AVTrack [31] introduces a novel single-stream tracking framework, featuring an adaptive ViT-based backbone and a prediction head. To improve the efficiency of the ViT, each transformer block in the ViT-based backbone, except for the first $n_{f}$ blocks, incorporates an Activation Module (AM). This module is trained to adaptively determine whether to activate the associated transformer block. As shown in Fig.3, the internal mechanism operates as follows: AM first slices the input tokens and passes them through a Linear layer followed by a Sigmoid function to compute the activation probability. For $i$-th ViT block ($i>n_{f}$), if its corresponding probability exceeds a predefined threshold, the block is activated; otherwise, it is deactivated. This mechanism enables adaptive, input-dependent computation and enhances the model’s efficiency. The framework takes a pair of images as input, comprising a template denoted as $Z\in\mathbb{R}^{3\times H_{z}\times W_{z}}$ and a search image denoted as $X\in\mathbb{R}^{3\times H_{x}\times W_{x}}$. These images are split into patches of size $P\times P$, and the number of patches for $Z$ and $X$ are $P_{z}=H_{z}\times W_{z}/P^{2}$ and $P_{x}=H_{x}\times W_{x}/P^{2}$, respectively. The features extracted from the backbone are fed into the prediction head to generate tracking results. To obtain view-invariant representation with ViTs, we maximize the mutual information (MI) between the feature representations of two different views of the target, i.e., the template image and the target patch in the search image. During the training phase, since the ground truth localization of the target in the search image is known, we can obtain the feature representation of the subsequent view from the representation of the search image using interpolation techniques.
Building on the proposed AVTrack, we introduce a simple yet effective MI maximization-based multi-teacher knowledge distillation (MD) framework to develop a more efficient UAV tracker, called AVTrack-MD. During the training of the AVTrack-MD model (i.e., the student model), the weights of AVTrack (i.e., the teacher model) remain fixed, while both teacher and student models receive the same inputs $Z$ and $X$. Let $\mathfrak{B}_{T}$ and $\mathfrak{B}_{S}$ represent the backbones of the teacher and student, respectively. In our implementation, $\mathfrak{B}_{S}$ shares the same structure as the $\mathfrak{B}_{T}$ but uses a smaller ViT layer. Feature-based knowledge distillation transfers knowledge from the teacher models’ backbone features to the student model by maximizing the MI between the averaged softened features of the multi-teacher models and the student model’s softened features. The details of these design principles will be elaborated in the subsequent subsections.
@@ III-B Activation Module (AM)
The Activation Module selectively activates transformer blocks based on input, enabling adaptive adjustments to the ViT architecture and ensuring that only the essential blocks are activated for effective tracking. To enhance efficiency, the AM processes a subset of tokens representing both the target template and the search image. Its output determines whether the current transformer block is activated. If not, the tokens from the preceding block are directly propagated to the next without further processing. This mechanism is consistently applied across all transformer blocks. Specifically, let’s consider the $i$-th layer ($i>n_{f}$). We denote the total number of tokens by $\mathcal{K}$, the embedding dimension of each token by $d$, and all the tokens output by the (i1)(i-1)-th layer by $\mathbf{t}_{1:\mathcal{K}}^{i-1}(Z,X)\in\mathbb{R}^{\mathcal{K}\times d}$. The slice of all tokens generated by the (i1)(i-1)-th transformer block is expressed as $\mathbf{e}_{1}^{\textup{T}}\mathbf{t}_{1:\mathcal{K}}^{i-1}(Z,X):=\textbf{r}^{i-1}\in\mathbb{R}^{\mathcal{K}}$, where $\mathbf{e}_{1}^{\textup{T}}=[1,0,...,0]\in\mathbb{R}^{\mathcal{K}}$ is a standard unit vector in $\mathbb{R}^{\mathcal{K}}$, which is used to extract the first token slice from all tokens output by the (i1)(i-1)-th transformer block. This design choice serves to isolate the specific token representation $r^{i-1}$ from the full token set $\mathbf{t}_{1:\mathcal{K}}^{i-1}(Z,X)$ before further processing via the linear layer $\mathfrak{L}^{i}$. This token carries global contextual information and serves as a compact representation for determining the activation state of the subsequent transformer block via the AM. This design ensures minimal overhead while retaining high-level semantics for adaptive computation. Formally, the Activation Module (AM) at layer $i$ is expressed as:
@@
where $p^{i}\in[0,1]$ represents the activation probability of the $i$-th transformer block, $\sigma(x)=1/(1+e^{-x})$ indicates the sigmoid function. The sigmoid function is versatile and finds applications in various machine learning tasks where non-linear transformations or probabilistic outputs are required. In our work, the activation mechanism in each transformer block is determined by the sigmoid function, which helps introduce non-linearity into the AM’s decision-making process. Specifically, the sigmoid function maps the output value of the linear layer to a range between 0 and 1, which represents the activation probability of the associate transformer block. If $p^{i}>\beta$, where $\beta\in(0.5,1)$ is the activation probability threshold, the transformer block at layer $i$ will be activated; otherwise, it is deactivated and the output tokens from the (i1)(i-1)-th layer will be fed into the $(i+1)$-th block directly. Let $\mathcal{N}$ denote the total number of transformer blocks in the given ViT. Theoretically, deactivating all $\mathcal{N}$ blocks simultaneously would result in no correlation being computed between the template and search image. To avoid such unfavorable conditions, the first $n_{f}$ layers are mandated to remain activated. This strategy helps alleviate computational burdens associated with AM, as these initial layers are typically essential, providing foundational information on which high-level and more abstract features and representations can be built. Another extreme case occurs when all transformer blocks are activated for any input, allowing the model to more efficiently minimize classification and regression losses, as larger models have greater fitting capacity. To address this, we introduce a block sparsity loss, $\mathcal{L}_{spar}$, which penalizes a higher mean probability across all adaptive layers, encouraging the deactivation of many blocks on average to improve efficiency. The block sparsity loss is defined as follows:
@@ III-C View-Invariant Representations (VIR) via Mutual Information Maximization
Figure 4: An overview of the proposed MI maximization-based multi-teacher knowledge distillation. Note that $\mathcal{F}_{\tau}$’ denotes the randomly shuffled of $\mathcal{F}_{\tau}$.
@@ III-D MI Maximization-based Multi-Teacher Knowledge Distillation (MD)
In distillation training, the student model is trained using a weighted sum of $\mathcal{L}_{MD}$ and the total loss function used for training the teacher model.
@@ III-E Prediction Head and Training Objective
For the tracking task, we employ the weighted focal loss [82] for classification and a combination of $L_{1}$ loss and GIoU loss [83] for bounding box regression. The total loss function for tracking prediction is defined as follows:
$\mathcal{L}_{pred}=\mathcal{L}_{cls}+\lambda_{iou}\mathcal{L}_{iou}+\lambda_{L_{1}}\mathcal{L}_{L1}$ (9)
where the constants $\lambda_{iou}=2$ and $\lambda_{L_{1}}=5$ are the same as in [21]. The AVTrack framework is trained end-to-end using the overall loss $\mathcal{L}_{overall}=\mathcal{L}_{pred}+\gamma\mathcal{L}_{spar}+\kappa\mathcal{L}_{vir}$, where $\gamma$ is set to 50 and $\kappa$ to 0.0001 after initialization with pretrained ViT weights. After this training, we apply the proposed MI maximization-based multi-teacher knowledge distillation framework to obtain a student model that better balances performance and efficiency. Specifically, during the distillation phase, the total loss $\mathcal{L}_{overall}^{*}=\mathcal{L}_{pred}+\eta\mathcal{L}_{MD}$ is employed for end-to-end distillation training, with the weight $\eta$ of $\mathcal{L}_{MD}$ is set to $0.1\times 10^{-3}$.
@@ IV Experiments
In this section, we evaluate our tracker on six publicly widely-used UAV tracking benchmarks, including DTB70 [84], UAVDT [85], VisDrone2018 [86], UAV123 [44], UAV123@10fps [44], and WebUAV-3M [42]. Specifically, we conduct a thorough comparison with over 38 existing state-of-the-art (SOTA) trackers, using their results obtained from running the official codes with the corresponding hyperparameters. Please note that our evaluation is performed on a PC that was equipped with an i9-10850K processor (3.6GHz), 16GB of RAM, and an NVIDIA TitanX GPU. To facilitate clearer comparison among different tracker types, we categorize them into two groups: lightweight trackers and deep trackers.
@@ IV-A Implementation Details
Model. The AVTrack [31] features three trackers for evaluation: AVTrack-ViT, AVTrack-DeiT, and AVTrack-EVA, with each employing a lightweight ViT as its backbone: ViT-tiny [101], DeiT-tiny [102], and EVA-tiny [103], respectively. Building on it, AVTrack-MD introduces the novel MI maximization-based multi-teacher knowledge distillation (MD) framework to develop three corresponding improved trackers: AVTrack-MD-ViT, AVTrack-MD-DeiT, and AVTrack-MD-EVA. The head and input sizes for AVTrack and AVTrack-MD are configured identically. The head consists of a stack of four Conv-BN-ReLU layers, with input sizes set to $256^{2}$ for the search image and $128^{2}$ for the template image.
@@ IV-B Comparison with Lightweight Trackers
1) Overall Performance: We compare our methods with 22 SOTA lightweight trackers across five renowned UAV tracking datasets. The evaluation results are presented in Table I. From table, our AVTrack-DeiT [31] outperforms all SOTA trackers except Aba-ViTrack [1] across these five benchmarks in terms of average (Avg.) precision (Prec.) and success rate (Succ.). Specifically, RACF [91] exhibits the best performance among DCF-based trackers, with Avg. Prec. and Succ. of 74.6% and 51.2%, respectively. Among CNN-based trackers, DRCI [95] stands out with the highest Avg. Prec. and Succ. of 79.8% and 59.1%. However, even the best methods among DCF- and CNN-based trackers fall short of the worst method among ViT-based trackers. All ViT-based trackers achieve Avg. Prec. and Succ. exceeding 80.0% and 62.0%, respectively, clearly surpassing the CNN-based methods and significantly outperforming the DCF-based ones. Regarding GPU speed, the top three trackers are our proposed methods: AVTrack-MD-EVA, AVTrack-MD-DeiT, and AVTrack-MD-ViT, achieving tracking speeds of 334.4 FPS, 310.6 FPS, and 303.1 FPS, respectively. In terms of CPU speed, all our methods deliver real-time performance on a single CPU [^1] It is important to note that the real-time performance discussed in this work is applicable only to platforms that are similar to or more advanced than ours.. Obviously, DCF-based trackers are the most efficient UAV trackers, as all methods achieving speeds above 80 FPS belong to this category. However, these fastest DCF-based trackers are unable to achieve performance comparable to our methods. For instance, the fastest KCF [5] tracker achieves an average success rate of only 31.6%, approximately half of the performance attained by our methods. Additionally, DCF-based methods often incur substantial costs to enhance tracking precision. Specifically, RACF exhibits the best performance among DCF-based trackers; however, it runs at only 35.6 FPS, which is significantly lower than the slowest CPU speed of our trackers. Although CNN-based trackers can achieve tracking speeds comparable to our ViT-based trackers, the latter significantly outperform the former in overall performance. Although Aba-ViTrack achieves the highest average precision of 85.3% and the highest average success rate of 64.7%., our AVTrack-DeiT secures second place with only slight gaps of 1.2% and 0.3%, respectively. Furthermore, our AVTrack-DeiT demonstrates higher speeds than Aba-ViTrack. Notably, AVTrack-DeiT runs up to 41%/18% faster than Aba-ViTrack in terms of GPU/CPU speed. When the proposed MI-based multi-teacher knowledge distillation framework (MD) is applied to AVTrack-*, the resulting student model, AVTrack-MD-*, demonstrates a smaller performance drop, with some even achieving performance improvements, all while delivering a significant speed boost. Specifically, AVTrack-MD-* models show varying trade-offs: AVTrack-MD-ViT achieves 0.5% and 0.9% gains in Avg. Prec. and Succ., respectively, with 21.1% GPU and 8.7% CPU speed improvements; AVTrack-MD-EVA gains 0.4% in precision and a minor 0.1% drop in success rate, while also improving GPU and CPU speeds by 17.9% and 6.8%; AVTrack-MD-DeiT has a slight performance drop of 0.5% in precision and 0.3% in success rate but still benefits from significant 20.9% GPU and 8.9% CPU speed gains. These results highlight the advantages of our method and justify its SOTA performance for UAV tracking.
@@
PRL-Track [96] 62.3 47.1 12.1 7.4 33.8
LiteTrack [97] 69.1 54.1 28.3 7.3 32.1
SMAT [98] 68.9 53.9 8.6 3.2 32.7
Aba-ViTrack [1] 70.4 55.3 8.0 2.4 37.3
ABDNet [94] 63.9 48.7 12.3 8.3 33.2
SGDViT [93] 61.3 45.7 23.3 11.3 31.7
TCTrack++ [92] 63.9 48.3 17.6 8.8 32.5
TCTrack [14] 61.9 45.7 8.5 6.9 34.4
SiamAPN [46] 62.5 45.1 14.5 7.9 35.2
HiFT [3] 60.9 45.5 9.9 7.2 35.6
@@
2) Performance on WebUAV-3M [42]: As illustrated in Table II, we have conducted a comparison of our AVTrack-DeiT and AVTrack-MD-DeiT with ten lightweight SOTA trackers on WebUAV-3M. Our AVTrack-DeiT surpasses all other lightweight trackers in success rate, achieving a 1.1% improvement over the second-place Aba-ViTrack while maintaining comparable precision with only a 0.4% difference. Notably, our AVTrack-MD-DeiT experiences less than a 1.0% drop in both precision and success rate, while achieving a 9.2% improvement in AGX speed. These results further underscore the effectiveness of our approach.
@@
3) Efficiency Comparison: To further demonstrate that our tracker achieves a better trade-off between accuracy and efficiency, we compare our methods against ten SOTA lightweight trackers, in terms of floating point operations (FLOPs), the number of parameters (Params.), and the speed (AGX.FPS) on the Nvidia Jetson AGX Xavier edge device. The results are shown in Table II. Notably, since our AVTrack-DeiT tracker features adaptive architectures, the FLOPs and parameters range from minimum to maximum values. As observed, both the minimum FLOPS and Params. of our AVTrack-DeiT are notably lower than those of all the SOTA trackers, and even its maximum values are lower than those of most SOTA trackers. Furthermore, our AVTrack-MD-DeiT achieves the lowest FLOPs (1.5G) and parameters (5.3M), except for AVTrack-DeiT, while also delivering the fastest speed at 46.1 AGX.FPS. This comparison in terms of computational complexity also underscores the efficiency of our methods.
TABLE III: Compwith state-of-the-art methods on two long-term UAV tracking benchmarks: UAV20L [44] and UAVTrack112L [43].
UAV20L UAVTrack112L
Tracker Prec. Succ. Prec. Succ.
AVTrack-DeiT 85.6 68.2 80.2 63.9
AVTrack-MD-DeiT 83.8 66.6 78.8 62.7
PRL-Track [96] 71.8 57.1 78.4 59.7
LiteTrack [97] 84.8 66.7 78.8 62.1
SMAT [98] 82.3 65.9 76.4 60.8
Aba-ViTrack [1] 85.2 67.9 81.1 64.2
ABDNet [94] 74.4 59.5 77.1 60.2
SGDViT [93] 69.7 55.8 73.8 55.1
TCTrack++ [92] 74.0 58.8 78.7 60.0
TCTrack [14] 75.1 59.6 72.0 53.7
SiamAPN[46] 70.9 55.6 78.0 57.8
HiFT [3] 78.0 56.8 70.8 53.1
4) Performance in Long-Term Tracking Scenarios: To better demonstrate the effectiveness of the proposed methods in long-term tracking, we have evaluated the performance of the proposed tracker on two long-term UAV tracking benchmarks: UAV20L [44] and UAVTrack112L [43], as shown in Table III. As observed, our AVTrack-DeiT achieves the best performance on UAV20L and the second-best on UAVTrack112L. Additionally, AVTrack-MD-DeiT performs comparably to AVTrack-DeiT and ranks third on UAVTrack112L. Although Aba-ViTrack achieves performance comparable to our trackers, as shown in Table I, its tracking speed is significantly slower than that of our methods. The results demonstrate that our tracker achieves competitive performance compared to the state-of-the-art methods, confirming its robustness in long-term tracking scenarios.